Show that potential function U= x2+y2+2z2 satisfies Laplace equation


1
Expert's answer
2021-02-17T14:43:27-0500

"\nabla^2U=\frac{\partial^2U}{\partial x^2}+\frac{\partial^2U}{\partial y^2}+\frac{\partial^2U}{\partial z^2}=0,"

"\frac{\partial U}{\partial x}=2x," "\frac{\partial U}{\partial y}=2y," "\frac{\partial U}{\partial z}=4z,"

"\frac{\partial^2U}{\partial x^2}=2," "\frac{\partial^2U}{\partial y^2}=2," "\frac{\partial^2U}{\partial z^2}=4,"

"\nabla^2U=2+2+4=8 \not =0,"

function U does not satisfy Laplace equation.